Your client requests you to evaluate two different pump systems for a facility. Asuume an interest rate of 9% per year and each system has a life span of 15 years.

System 1: Initial cost = $55,000 O&M cost = $2500 for the first year and increasing $100 per year for the remainder of the life span Salvage Value = $15,000

System 2: Initial cost = $50,000 O&M cost = $2000 for the first year and increasing $150 per year for the remainder of the life span Salvage Value = $16,000

Determine the most cost effective solution for your client by evaluating the present worth.
Answer #1

Net present cost of System 1 = 55000 + 2500 * (P/A,9%,15) + 100 * (P/G,9%,15) - 15000 * (P/F,9%,15)

= 55000 + 2500 *8.060688 + 100 *43.806865 - 15000 *0.274538

= 75414.34

Net present cost of System 2 = 50000 + 2000 * (P/A,9%,15) + 150 * (P/G,9%,15) - 16000 * (P/F,9%,15)

= 50000 + 2000 *8.060688 + 150 *43.806865 - 16000 *0.274538

= 68299.80

As Net present cost of System 2 is lower it should be selected